OpenStudy (babynini):

Vectors U = (-6,7) find 2u I'm not sure why this answer isn't coming out right o.o

OpenStudy (babynini):

@IrishBoy123

OpenStudy (loser66):

2*(-6 ) =? 2*7 =?

OpenStudy (babynini):

12, 14

OpenStudy (loser66):

14 is correct , but 12 is not

OpenStudy (babynini):

um..-12?

OpenStudy (loser66):

yes, dat sit

OpenStudy (babynini):

Why don't the negatives cancel?

OpenStudy (loser66):

2u means 2*u not u^2 which is u*u

OpenStudy (loser66):

for example, if the number is -3, then 2*(-3) = -6, while (-3)^2 = (-3)*(-3) = 9 . DIFFERENT!!!

OpenStudy (babynini):

hrm hrm ok

OpenStudy (babynini):

thanks!